OpenStudy (anonymous):

What is the equation of a line passing through (3, 4) and having a slope of -5?

OpenStudy (anonymous):

Use the point slope formula: y - y1 = m(x - x1) where m is the slope and (x1,y1) is any point on the line y - 4 = -5(x - 3)

OpenStudy (anonymous):

A. y = 5x + 19 B. y = -5x + 11 C. y = -5x + 19D. .

OpenStudy (anonymous):

?

OpenStudy (anonymous):

A Just simply what I wrote and isolate y on the left side of the equation You get answer A